can 2,3, and 4 be the sides of a triangle
Found 2 solutions by fcabanski, Alan3354:
Answer by fcabanski(1391)

The requirement for the side lengths of triangles is that the sum of any two sides has to be larger than the third side.


2+3 = 5 > 4


4+2 = 6 > 3


3+4 = 7 > 2


The sides of a triangle can be 2, 3 and 4.


Answer by Alan3354(69443)

can 2,3, and 4 be the sides of a triangle
-------------
Yea, why not?